# Universal Time-Series Representation Learning: A Survey

### v1

@@ 1.4. Contributions
This paper aims to identify what are the essential elements in designing state-of-the-art representation learning methods for time series and how these elements affect the quality of the learned representations. To the best of our knowledge, this is the first survey on universal time-series representation learning. We propose a novel taxonomy for learning universal representations of time series from the novelty perspectives—whether the main contribution of a paper focuses on what part of the design elements discussed above—to summarize the selected studies. Table 2 summarizes and compares the reviewed articles based on the proposed taxonomy. From the perspective of neural architectures, we summarize the changes made on both the module level and network level of neural architectures used for representation learning. From the learning perspective, we classify what objectives are used to make the learned representations generalizable to various downstream tasks. Last, we also categorize data-centric methods for the papers that focus particularly on improving the quality of the training data. Overall, our main contributions are as follows.
  • We conduct an extensive literature review of universal time-series representation learning based on a novel and up-to-date taxonomy that categorizes the reviewed methods into three main categories: neural architectural, learning-focused, and data-centric approaches.
@@
1.5. Article Organization
Section 2 introduces the definitions and specific background knowledge regarding time-series representation learning. In Section 3, we present a comprehensive review of the methods focusing on neural architecture aspects. We then discuss the methods focusing on deriving new learning objectives in Section 4. Section 5 gives a review on data-centric approaches. In addition, we discuss the evaluation protocol for time-series representation learning and promising future research directions in Section 6 and Section 7, respectively. Finally, Section 8 concludes this survey.
Table 2. Taxonomy of universal representation learning for time series.
Design Element Category Section Technique Key References
Neural Architectures (Neural Architectural Approaches) Basic Block Combination 3.1 Network-Level Combination (Chen et al., 2021b; Liu et al., 2020; Zerveas et al., 2021; Jhin et al., 2021; Abushaqra et al., 2022; Jhin et al., 2022; Kim et al., 2023a; Rubanova et al., 2019b; Sanchez et al., 2019; Xie et al., 2022; Han et al., 2020; Choi and Kang, 2023; Rahman et al., 2021; Franceschi et al., 2019; Wang et al., 2018; Yang and Hong, 2022; Li et al., 2023; Liang et al., 2023a; Fraikin et al., 2023)
Module-Level Combination (Anand and Nayak, 2021; Nguyen et al., 2023; Wu et al., 2023; Chen et al., 2019; Zhou et al., 2023; Luetto et al., 2023; Guo et al., 2021)
Innovative Redesign 3.2 Network-Level Design (Liang et al., 2023b; Zhang et al., 2023b; Bianchi et al., 2019; Schirmer et al., 2022; Biloš et al., 2023; Ansari et al., 2023; Wu et al., 2018; Sun et al., 2021; Liu et al., 2019; Fons et al., 2022; Wang et al., 2023a; Li et al., 2022; Cai et al., 2021; Lee et al., 2022a; Wang et al., 2023b; Chen et al., 2023; Zhang et al., 2023c; Zhong et al., 2023; Xu et al., 2023)
Module-Level Design (Ma et al., 2019; Wang et al., 2023c; Gorbett et al., 2023; Shukla and Marlin, 2021; Sener et al., 2020; Chowdhury et al., 2022; Liu et al., 2022; Naour et al., 2023; Lin et al., 2023)
Learning Objectives (Learning-Focused Approaches) Supervised Learning 4.1 Task-Specific Loss (Ma et al., 2019; Rahman et al., 2021; Hadji et al., 2021; Zhong et al., 2023)
Unsupervised Learning 4.2 Masked Prediction (Zerveas et al., 2021; Luetto et al., 2023; Chowdhury et al., 2022)
Input Reconstruction (Sanchez et al., 2019; Yuan et al., 2019; Tonekaboni et al., 2022; Nguyen et al., 2018; Dong et al., 2023; Li et al., 2023; Chen et al., 2019)
Self-Supervised Learning 4.3 Non-Contrasting Loss (Guo et al., 2022; Haresh et al., 2021; Liang et al., 2022; Wang et al., 2020, 2021; Duan et al., 2022; Fang et al., 2023; Fraikin et al., 2023)
Contrasting Loss (Liang et al., 2023b; Zhang et al., 2022a; Kong et al., 2020; Tonekaboni et al., 2022; Ding et al., 2022; Kim et al., 2023a; Chen et al., 2022; Zhang and Crandall, 2022; Haresh et al., 2021; Qing et al., 2022; Morgado et al., 2020; Nguyen et al., 2023; Han et al., 2020; Zhang et al., 2023d; Choi and Kang, 2023; Qian et al., 2022; Chowdhury et al., 2023; Chen et al., 2021c; Zhang et al., 2022b; Hajimoradlou et al., 2022; Qian et al., 2021; Liu et al., 2022; Dave et al., 2022; Yang et al., 2023a; Jenni and Jin, 2021; Jiao et al., 2020; Yang et al., 2022a; Somaiya et al., 2022; Yue et al., 2022; Tonekaboni et al., 2021; Franceschi et al., 2019; Lee et al., 2022a; Shin et al., 2023; Wang et al., 2023b; Liu et al., 2023; Zhang et al., 2023e; Sun et al., 2023; Eldele et al., 2021)
Training Data (Data-Centric Approaches) Data Augmentation 5.1 Random Augmentation (Yue et al., 2022; Zhang et al., 2022b, 2023e)
Policy-based Augmentation (Aboussalah et al., 2023; Kim et al., 2023b; Luo et al., 2023; Kim et al., 2023c; Zhang and Crandall, 2022; Yang and Hong, 2022; Chen et al., 2022; Shin et al., 2023; Yang et al., 2022a; Demirel and Holz, 2023)
Decomposition and Transformation 5.2 Time-Series Decomposition (Zeng et al., 2021; Behrmann et al., 2021; Wang et al., 2018; Yang et al., 2022b; Fang et al., 2023)
Input Space Transformation (Anand and Nayak, 2021; Biloš et al., 2022; Lee et al., 2022b; Wu et al., 2023; Chanpuriya et al., 2023; Zhong et al., 2023; Xu et al., 2023)
Methods for Sample Selection 5.3 Generative Methods (Nguyen et al., 2023; Zhao et al., 2023)
(Re-)Sampling Strategies (Chen et al., 2021b; Franceschi et al., 2019)
@@ 2. Preliminaries
This section presents definitions and notations used throughout this paper, descriptions of downstream tasks, unique properties in time series, and basic building blocks of neural architectures.
@@ 2.1. Definitions
An irregularly-sampled time series is a time series that the intervals between observations are not consistent or regularly spaced. Thus, the time intervals between $(\mathbf{x}_{1},\mathbf{x}_{2})$ and $(\mathbf{x}_{2},\mathbf{x}_{3})$ is unequal, as illustrated in Fig. 4. It is often encountered in situations where data is collected opportunistically or when events occur irregularly and sporadically, e.g., sensor malfunctions, leading to varying time gaps between observations.
@@
A crucial measure to assess the quality of a representation learning method, i.e., the encoder $f_{e}$, is its ability to produce the representations $\mathbf{Z}$ that effectively facilitate downstream tasks, either with or without fine-tuning (see Section 6). Once we obtain the latent representation $\mathbf{Z}$, we can use it as input for downstream tasks to evaluate its actual performance. Here, we define the most common downstream tasks as follows.
@@
(a) Regularly Sampled Time Series.
(b) Irregularly Sampled Time Series.
@@
Note that the above definitions are generally based on the raw time series $\mathbf{X}$. Following Definition 2.3, we can use the corresponding representation $\mathbf{Z}=f_{e}(\mathbf{X})$ to perform the above downstream tasks instead of the raw data.
@@ 2.2.1. Temporal Dependency
Time series exhibits dependencies on the time variable, where a data point at a given time correlates with its previous values. Given an input $\mathbf{x}_{t}$ at time $t$, the model predicts $y_{t}$, but the same input at a later time could be a different prediction. Therefore, windows or subsequences of past observations are usually included as inputs to the model to learn such temporal dependency. The length of windows for capturing the time dependencies could also be unknown. In addition, there are local and global temporal dependencies. The former is usually associated with abrupt changes or noises, while the latter is associated with collective trends or recurrent patterns.
@@ 2.3.5. Attention-based Networks
2.3.6. Neural Ordinary Differential Equations (Neural ODE)

### v3

@@ 1.3. Contributions
This paper aims to identify what are the essential elements in designing state-of-the-art representation learning methods for time series and how these elements affect the quality of the learned representations. To the best of our knowledge, this is the first survey on universal time-series representation learning. We propose a novel taxonomy for learning universal representations of time series from the novelty perspectives—whether the main contribution of a paper focuses on what part of the design elements discussed above—to summarize the selected studies. Table 2 outlines and compares the reviewed articles based on the proposed taxonomy. We begin by exploring the essence of data-driven methods, categorizing data-centric approaches from the papers that primarily aim to enhance the usefulness of the training data. From the perspective of network architectures, we review the evolution of neural architectures by focusing on how these advancements enhance the quality of learned representations. From the learning perspective, we classify how different objectives enhance the generalizability of learned representations across diverse downstream tasks. Overall, our main contributions are as follows.
  • We conduct an extensive literature review of universal time-series representation learning based on a novel and up-to-date taxonomy that categorizes the reviewed methods into three main categories: data-centric, neural architectural, and learning-focused approaches.
@@
Article Organization. Section 2 introduces the definitions and specific background knowledge regarding time-series representation learning. Section 3 gives a review on data-centric approaches. In Section 4, we present an extensive review of the methods focusing on neural architecture aspects. We then discuss the methods focusing on deriving new learning objectives in Section 5. In addition, we discuss the evaluation protocol for time-series representation learning and promising future research directions in Section 6 and Section 7, respectively. Finally, Section 8 concludes this survey.
[…]
Illustrations of regularly and irregularly sampled multivariate time series with V = 3.
@@ 2. Preliminaries
This section presents definitions and notations used throughout this paper, descriptions of downstream time-series analysis tasks, and unique properties of time series.
@@ 2.1. Definitions
An irregularly-sampled time series (ISTS) is a time series where the intervals between observations are not consistent or regularly spaced. Thus, the time intervals between $(\mathbf{x}_{1},\mathbf{x}_{2})$ and $(\mathbf{x}_{2},\mathbf{x}_{3})$ is unequal, as illustrated in Fig. 4. It is often encountered in situations where data is collected opportunistically or when events occur irregularly and sporadically, e.g., sensor malfunctions, leading to varying time gaps between observations.
@@
A crucial measure to assess the quality of a representation learning method, i.e., the encoder $f_{e}$, is its ability to produce the representations $\mathbf{Z}$ that effectively facilitate downstream tasks, either with or without fine-tuning (see Section 6). Once we obtain $\mathbf{Z}$, we can use it as input for downstream tasks to evaluate its actual performance. Here, we define the common downstream tasks as follows.
@@
Following Definition 2.3, we can use the corresponding representation $\mathbf{Z}=f_{e}(\mathbf{X})$ to perform the above downstream tasks instead of the raw time series $\mathbf{X}$.
@@ 2.2.1. Temporal Dependency
Time series exhibits dependencies on the time variable, where a data point at a given time correlates with its previous values. Given an input $\mathbf{x}_{t}$ at time $t$, the model predicts $y_{t}$, but the same input at a later time could be a different prediction. Therefore, windows or subsequences of past observations are usually included as inputs to the model to learn such temporal dependency. The length of windows for capturing the time dependencies could also be unknown. There are local and global temporal dependencies. The former is usually associated with abrupt changes or noises, while the latter is associated with collective trends or recurrent patterns.
@@ 2.3.5. Attention-based Networks
2.3.6. Neural Differential Equations (NDE)